Mission Plan – April 11th

Crew 245 Mission Plan 11th April 2022

SOL: 01
Name of person filing report: Vittorio Netti

The SMOPS crew has 13 experiments, both passive and active, mainly focused on crew monitoring and support. They are divided into four main categories: crew health, space suits and monitoring, space support technology, and planetary science. For the first category, Benjamin Pothier (crew journalist) will make use of electroencephalography to observe the changes in his brain during meditation sessions and we will perform a spit test for cortisol measurement to estimate two crew members’ stress levels before and after an EVA (PI: Nadia Maarouf, HSO). In the second category, crew gear (prototype flight suit and boots) and wearable sensors will be tested to monitor the movements of crew members during EVAs (PI: Paolo Guardabasso, XO), and each crew member will also wear an undergarment for continuous monitoring (PI: Nadia). We will also test some technologies that will support astronaut activities in future missions, such as purification of air from bacteria and satellite communication (PI: Luca Rossettini, crew engineer), 3D scanning of station modules and geological features (PI: Benjamin), 3D Printing of tools and drone flight (PI: Vittorio Netti, commander). Last but not least, we’ll take advantage of the geological analogy between MDRS site and Mars, collecting samples in search of micrometeorites and DNA traces (PI: Simone Paternostro, crew scientist).

This challenging list of experiments will require intensive EVAs, with the following main objectives: 3D Scan of MDRS modules and of geological features around the MDRS site; test of 3D printed objects in sample collection and other operations; drone flight to search for interesting geological features; test of shoes, flight suits, lower suit, helmet for ergonomy on different terrains and record data with wearables; sample collection (10 to 30 locations) for micrometeoroids and DNA samples. 1 EVA walking in the area around Hab (4 locs), 1 South (4 locs), 1 East (4 locs), 1 West (4-8 locs); antenna assembly and maintenance.

Throughout the mission we will also perform outreach and filming activities: Benjamin is currently working on a documentary on analogue missions and he will film the crew performing operations with cameras and drones; there will be an attempt to leave a camera outside during an EVA and leave it on overnight for a sky timelapse. We will also take time to acknowledge our sponsors.

Commander, out

Commander Report – April 11th

many events to report for our report of Sol 1.

Due to the lack of information about the last three packages, we decided this morning to wait one last day "out-of-sim" in case a trip to Salt lake city is in order to pick them up. Luckily the situation was resolved around lunchtime, with the confirmation that private couriers have picked up the packages and will deliver them tomorrow eventually.

We spent the day planning and preparing the experiments for the next days and fixing different small hardware failures that affect the habitat and the mission tools. That included the water sensor, the batteries of three different suits, and one of the steps of the stair to the upper deck.

We had our first amazing meal cooked by Crew Journalist Benjamin Pothier, we are thinking to elect him the official SMOPS mission chef.

In the afternoon we also performed a full 3D scan of the RAM module, to test our advanced handheld lidar 3D scanner.

High winds are still hitting the habitat pretty hard. The external airlock door has been slammed open destroying the left wooden barrier. We have been able to close and secure back the door, but the situation was too dangerous to attempt repairs of the barrier.

The simulation will officially start tomorrow morning at 7.30 am. Two crew members, (XO Paolo Guardabasso and Crew Eng Luca Rossettini) will need to break the sim for about 1 hour to recover the last shipped objects as soon as they are delivered.

Commander out,

Vittorio Netti

Operations Report – April 11th

Crew 245 Operations Report 11-04-2022
SOL: 1
Name of the person filing report: Luca Rossettini

Non-nominal systems: Suits, Heating system, main door, side bar of the
Notes on non-nominal systems:
Water alarm is functional to be tested.
Suits: we request authorization to substitute the three batteries.
Where should we put disposable batteries (AA, AAA)?
The heating system is not working. We are using the heater from the
RAM at the upper deck.
Spirit rover used: No
Opportunity rover used: No
Curiosity rover used: No
Perseverance rover used: No
General notes and comments: None
Summary of Hab operations:
Static tank was refilled by Shannon almost to top
Water (static tank): 430 gallons
Water (loft tank): 11 gallons
Water Meter: 157452,5
Static to Loft Pump used: Yes
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: No
Summary of internet: nothing to report
Summary of suits and radios:
Radios charging.
We disconnected suit 3 from charging and the fan was turning on only
at maximum power. We checked the battery and the voltage was 6 V.
Same problem with suits 4 and 6 (they wouldn’t even turn on the
fan), that both had 4V in fully charged state.
We tried charging them with another charger (to see if that might be
the problem, but the low voltage did not change).
Suit 1, once disconnected, is working only at maximum speed.
Slightly reducing the speed causes the suit not to working.
We request authorization to substitute the three batteries.
Summary of GreenHab operations:
WATER USE: 0 gallons
Heater: Off
Supplemental light: Off
Summary of ScienceDome operations:
Checked inventory
Dual split (Heat/AC): Off
Summary of RAM operations: Not used, nothing to report
Summary of any observatory issues: Not used, nothing to report
Summary of health and safety issues:
The main door is now damaged due to high winds and cannot stay
closed. We used the red wired to secure it within the habitat, passing
through the airlock and being wired to a metal tube. This is preventing the
Martians to use the airlock and reduce the exits.
Furthermore, the side bar right outside the main door has been
damaged by the door slamming wide open into it. Tomorrow we will check if it
will be sufficient to screw it firmly or if other type of repair will be
Questions, concerns and requests to Mission Support:
Suits: we request authorization to substitute the three batteries.
Where should we put disposable batteries (AA, AAA)?
Concern: with the high winds the main door is not safe, it could fly
away or anyway get permanently damaged.

Sol Summary – April 11th

Crew 245 Sol Summary Report 11-04-2022
SOL: 1
Summary Title: A windy situation
Author’s name: Paolo Guardabasso

Mission status: Sim has been delayed all day waiting for favourable conditions for the unsuited rover training that did not yet happen. We decided to start sim tomorrow morning.

Sol Activity Summary: The crew performed EVA suit training and tested the SMOPS flight suits and boots. The first meal at the station was prepared (no insects were hurt in the process). The afternoon was dedicated to some repairs, checking EVA suits, inventory and experiment preparation. Because of the high wind, it was not possible to go out and perform the rover training. Around the start of CapCom the front door was opened by the wind and slammed, breaking the wooden rail of the stairs.

Look Ahead Plan: Later we might attempt to repair the stairs, or more likely in a future EVA. Tomorrow morning we will officially start sim in the morning. Some cargo is still on route, hence 2 members will have to break sim during the day to retrieve it in Hanksville. Hopefully we’ll perform 2 EVAs in the day.

Anomalies in work: High wind delayed the taking of a group picture (also that with the banner) and rover training.

Weather: High winds, average speed 45 kmh and with a maximum of 75 kmh.

Crew Physical Status: All nominal.

EVA: No suited EVAs performed.

Reports to be filled:
– Mission plan
– HSO report
– Operation report
– Journalist report
– Picture of the day
– Some pictures
– EVA request (2)

Support Requested: We need the Yuri’s banner to take a picture tomorrow in EVA.

Copyright © The Mars Society. All rights reserved. | Main Site